# Optimization in quantum metrology and entanglement theory using semidefinite programming

### v3

@@ Abstract
We discuss efficient methods to optimize the metrological performance over local Hamiltonians in a bipartite quantum system. For a given quantum state, our methods find the best local Hamiltonian for which the state outperforms separable states the most from the point of view of quantum metrology. We show that this problem can be reduced to maximize the quantum Fisher information over a certain set of Hamiltonians. We present the quantum Fisher information in a bilinear form and maximize it by an iterative see-saw (ISS) method, in which each step is based on semidefinite programming. We also solve the problem with the method of moments that works very well for smaller systems. We consider a number of other problems in quantum information theory that can be solved in a similar manner. For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion. Our approach is one of the efficient methods that can be applied for an optimization of the unitary dynamics in quantum metrology, the other methods being, for example, machine learning, variational quantum circuits, or neural networks. The advantage of our method is the fast and robust convergence due to the simple mathematical structure of the approach.
@@ I Introduction
Quantum entanglement is at the heart of quantum physics and several quantum information processing applications [1, 2, 3, 4]. It also plays a central role in quantum metrology. It is known that interparticle entanglement is needed to overcome the shot-noise limit of the precision in parameter estimation in a linear interferometer [5]. It has also been shown that the larger the achieved precision in parameter estimation with a multiparticle system, the larger the depth of entanglement the state must possess [6, 7]. However, not all entangled states are useful for metrology [8].
@@
The problem of the metrological performance of a bipartite state of qudits with a dimension larger than two is more complicated. In this case, the local Hamiltonians cannot be transformed into each other by local unitaries, and the best metrological performance achievable by separable states depends on the Hamiltonian. Thus, we introduce the metrological gain as a general way to characterize the quantum advantage of a system with arbitrary dimensions. It is defined as the ratio of the quantum Fisher information and the maximum of the quantum Fisher information over separable states maximized over local Hamiltonians [22]. The metrological gain is convex in the state, and it cannot decrease if additional copies or ancilla systems are added [22]. The optimization over the local Hamiltonian can also be carried out via an ISS method [22, 12, 23]. Thus, our method uses an ISS algorithm for optimizing the unitary dynamics rather than for optimizing the state. Note that instead of using the metrological gain, one can also consider the difference of the quantum Fisher information and its maximum for separable states [24], or even consider the difference with respect to a bound that depends on the state [25, 26, 27, 28].
@@
In this paper, we examine the optimization problem above in a great depth. We elaborate the point that local Hamiltonians with meaningful constraints form a convex set. Thus, we have to maximize the metrological gain over a convex set of local Hamiltonians. We find several algorithms to maximize the metrological gain. The first method is based on maximizing a quadratic expression using a see-saw over a bilinear expression, and its computational costs are minimal. We present other algorithms that need quadratic programming with quadratic constraints and relaxations. Such methods work for small systems. We test the methods and compare them to each other. We list a number of similar problems that can be solved in an analogous way, such as maximization of norms over a convex set.
@@
Our work is also related to recent efforts to solve various problems of numerical optimization in quantum metrology, including the combination of the tasks mentioned above, such as using an ISS for optimizing the quantum state and the unitary dynamics [22]. An ISS-based method has been used to optimize over a general dynamics using its Choi-Jamiolkowski representation [29]. An ISS-based method has also been used to optimize over adaptive strategies, when coherently probing several independent quantum channels [30]. An optimization over the probe state and the dynamics has been realized experimentally in a programmable quantum sensor realized in trapped cold ions [31]. For the identification of the optimal quantum metrological protocol in large systems, matrix product operators have been used [11]. Optimization methods, including ISS, has been considered in estimating spectral density of a stochastic signal field [32]. Finally, numerical optimization in quantum metrology in various complex scenarios has been considered including multiparameter estimation [33, 34, 35, 36, 37, 38], and there are even methods that are based on semidefinite programming [39].
@@
Our paper is organized as follows. In Sec. II, we review the basic notions of entanglement theory and quantum metrology. In Sec. III we discuss, how our optimization problem can be viewed as maximization over a convex set of Hamiltonians. We review the notion of metrological gain. In Sec. IV, we present various methods to optimize the quantum Fisher information over a local Hamiltonian. In Sec. V, we consider optimization problems in quantum information that can be solved in a similar way. For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion for entanglement detection [49, 50].
@@ II.1 Entanglement theory
In this section, we summarize basic notions of entanglement theory [1, 2]. A bipartite quantum state is called separable if it can be written as mixture of product states as [51]
@@ II.2 Quantum metrology
Here is the Hamiltonian of the evolution and is a parameter. Then, the metrological precision in estimating is bounded by the Cramér-Rao bound [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]
@@
the constant coefficients depend only on the eigenvalues of the probe state and are given as
whenever the denominator is not zero, and the matrix element of the Hamiltonian in the basis of the eigenvectors of the density matrix is given as
@@
where the superscripts "r" and "i" denote the real and imaginary parts, respectively.
@@ III Basic notions for optimizing the metrological performance
First, not all Hamiltonians are easy to realize. In particular, local Hamiltonians are much easier to realize than Hamiltonians with an interaction term, and such Hamiltonians appear most typically in quantum metrology. Thus, we will restrict the optimization for such Hamiltonians given as
@@
holds, where is some constant. Thus, it is easy to improve the metrological performance of any Hamiltonian by multiplying it by a constant factor. In order to obtain a meaningful optimization problem, we have to consider a formulation of such a problem where such a "trick" cannot be used to increase the quantum Fisher information to infinity. One option could be to divide the quantum Fisher information with the square of some norm or seminorm of the Hamiltonian, for which
@@
In this paper, we argue that it is advantageous to compare the metrological performance of a state to the best performance of separable states. Therefore, we will divide the quantum Fisher information of the state by the maximum quantum Fisher information achievable by separable states given as for some as [70, 10]
where and denotes the maximal and minimal eigenvalues, respectively.
@@ III.1 Seminorm for local Hamiltonians
Proof. We will now follows steps similar to those in the proof of Observation 1, showing that the three properties of seminorms are satisfied.
@@ III.2 Metrological gain
We remark that there are other quantities to characterize the metrological performance. One can consider difference of the quantum Fisher information and the its maximum for separable states, and include the possibility of non-unitary dynamics from the point of view of resource theories. This way, all entangled states turn out to be a useful resource [24]. One can also consider the difference the quantum Fisher information and a bound that depends on the state, where the difference is positive only for entangled states [25, 26, 27, 28].
@@ IV.1 See-saw optimization based on a bilinear form
where contains the ’s and contains the ’s and the ’s. The matrix is a positive semidefinite diagonal matrix. Minimizing such an expression over is a simple problem. On the other hand, maximizing it is difficult. A usual way of maximizing such an expression is as follows [76, 77, 78]. We replace the optimization over by a simultaneous optimization of the vectors and of a bilinear form as
@@
Based on this, we can write the following.
@@
Based on Eq. (53), we can carry out the following algorithm. Initially, we choose randomly. Then, we maximize over the matrix while keeping fixed. Such an optimization can be carried out, for instance, by semidefinite programming [79] or by simple matrix algebraic calculations, see Appendix B. Then, we maximize over while keeping fixed. Then, we optimize again over , etc. This process is illustrated in Fig. 3. Note that we do not have to verify that the random initial Hamiltonian fulfills the constraints Eq. (47), since a random initial Hamiltonian and for any as initial Hamiltonian will lead to the same optimal (An alternative formulation of the optimization problem is in Appendix C. With similar methods, we also analyze the optimization over a quantum state rather than over the Hamiltonian Appendix D.)
@@
Since is convex in and if we maximize it over a convex set of Hamiltonians, then it will take the maximum at the boundary of the set. This optimization is a hard task, as we have discussed. We have to start from several random initial conditions, and only some of them will find a solution.
@@
where is the maximally entangled state and is the fraction of the white noise. In particular, we considered and . For such states, the optimal Hamiltonian is known analytically [22]. We also considered two copies of the state and given as where the bipartite system consists of the parties and We also considered some bound entangled states that have been examined in the literature, and are often used to test various algorithms. Such states have a nontrivial structure, are highly mixed, but can be metrologically useful. We looked at the state based on an Unextendible Product Basis (UPB) [80, 81], the Horodecki state for [55], the metrologically useful bound entangled state in Ref. [10], and the private state given in Ref. [82, 12]. The results are shown in Table 1. The success probability of the method of Ref. [22] is similar, however, the formulation used in this paper makes it possible to design methods that provably find the optimum 111We started from a random initial Hermitian Hamiltonian, where the real and imaginary parts of the elements had a normal distribution with a zero mean and a unit variance, and they were independent from each other, apart from the requirement of the Hermitianity..
@@ IV.2 Quadratic programming with quadratic constraints
In the previous section, we considered methods that aim to maximize the quantum Fisher information over the Hamiltonian numerically. In practice, they are very efficient. Still, it is not certain that they find the global optimum and they might find a smaller value. In this section, we present methods that always find the global optimum. However, they work only for small systems. We also present simple approximations of such methods called, relaxation. They find either the maximum or a value larger than that. The method given in Sec. IV.1 looking for the maximum with the see-saw and the methods discussed in this section can be used together. The gap between the two results give information about the location of the maximum. If the two methods find the same value, then it is the global maximum.
The basic idea of the method of moments can be understood as follows. According to the Shor relaxation [89, *[][.OriginallypublishedinSov.J.Comput.Syst.Sci.25, 1-11(1987).]Shor1987Quadratic], the maximization of Eq. (51) is replaced by the maximization of
@@
Note that Eq. (57) can easily be coded in semidefinite programming. We can also add further and further conditions on that lead to lower and lower values for the maximum. Similar relaxation techniques have been used in entanglement theory to detect entanglement, non-locality [91], to obtain the maximum of an operator for separable states [92] or even to determine the ground state energy in spin one-half Heisenberg models [93].
@@
which is equivalent to Eq. (58) for all values. Thus, we do not need to carry out calculations for a range of values. A somewhat weaker constraint, which is computationally simpler is
We tested our methods for the same states we considered in Sec. IV.1. For the example, on a laptop computer, we could carry out calculations at level and Even level gave the exact result. For the examples, we could carry out calculations at level and and level gave the correct result. For the and examples, the level calculation gave already the correct result. The results are shown in Table 1.
@@ V.2 Maximization of the Wigner-Yanase skew information
Maximizing the Wigner-Yanase skew information can be used to lower bound the quantum Fisher information based on
@@ V.3 Searching for the eigenstate with maximal/minimal eigenvalue
In this section, we discuss how to look for the largest eigenvalues of a matrix with a see-saw method similar to the one discussed in Sec. IV.1. Interestingly, the see-saw is guaranteed to converge in this case.
@@
where is a Hermitian matrix and the right-hand side of Eq. (66) can immediately be used to define a see-saw algorithm. We shall show that such an algorithm is equivalent to the von Mises method (power iteration) for finding the maximal eigenvalue of a positive self-adjoint operator.
Before presenting the see-saw, let us consider the Cauchy-Schwarz inequality
@@
Now, the see-saw procedure is as follows:
@@
We can now recognize the steps of the power iteration, which is known to converge as long as the starting vector ( in our case) overlaps with the eigenvector corresponding to the maximal eigenvalue of .

### v4

# Iterative optimization in quantum metrology and entanglement theory using semidefinite programming

@@ Abstract
We discuss efficient methods to optimize the metrological performance over local Hamiltonians in a bipartite quantum system. For a given quantum state, our methods find the best local Hamiltonian for which the state outperforms separable states the most from the point of view of quantum metrology. We show that this problem can be reduced to maximizing the quantum Fisher information over a certain set of Hamiltonians. We present the quantum Fisher information in a bilinear form and maximize it by an iterative see-saw (ISS) method, in which each step is based on semidefinite programming. We also solve the problem with the method of moments that works very well for smaller systems. Our approach is one of the efficient methods that can be applied for an optimization of the unitary dynamics in quantum metrology, the other methods being, for example, machine learning, variational quantum circuits, or neural networks. The advantage of our method is the fast and robust convergence due to the simple mathematical structure of the approach. We also consider a number of other problems in quantum information theory that can be solved in a similar manner. For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion.
@@ I Introduction
Quantum entanglement is at the heart of quantum physics and several quantum information processing applications [1, 2, 3, 4]. It also plays a central role in quantum metrology. It is known that interparticle entanglement is needed to overcome the shot-noise limit of the precision in parameter estimation in a linear interferometer [5]. It has also been shown that the larger the achieved precision in parameter estimation with a multiparticle system, the larger the depth of entanglement the state must possess [6, 7]. However, not all entangled states are useful for metrology [8]. Even bound entangled states, which are considered very weakly entangled, can be more useful for metrology than some distillable entangled states [9, 10, 11].
@@
The problem of the metrological performance of a bipartite state of qudits with a dimension larger than two is more complicated. In this case, the local Hamiltonians cannot be transformed into each other by local unitaries, and the best metrological performance achievable by separable states depends on the Hamiltonian. Thus, we introduce the metrological gain as a general way to characterize the quantum advantage of a system with arbitrary dimensions. It is defined as the ratio of the quantum Fisher information and the maximum of the quantum Fisher information over separable states maximized over local Hamiltonians [23]. The metrological gain is convex in the state, and it cannot decrease if additional copies or ancilla systems are added [23]. In the multiparticle case, the metrological gain is related to multiparticle entanglement, as it lower bounds the entanglement depth [24]. The optimization over the local Hamiltonians can also be carried out via an ISS method [23, 11, 24]. Thus, our method uses an ISS algorithm for optimizing the unitary dynamics rather than for optimizing the state. Note that instead of using the metrological gain, one can also consider the difference of the quantum Fisher information and its maximum for separable states [25], or even consider the difference with respect to a bound that depends on the state [26, 27, 28, 29].
@@
In this paper, we examine the optimization problem above in a great depth. We elaborate the point that local Hamiltonians with meaningful constraints form a convex set. Thus, we have to maximize the metrological gain over a convex set of local Hamiltonians. We find several algorithms to maximize the metrological gain. The first method is based on maximizing a quadratic expression using a see-saw over a bilinear expression, and its computational costs are minimal. We present other algorithms that need quadratic programming with quadratic constraints and relaxations. Such methods work for small systems. We test the methods and compare them to each other. In the second half of the paper, we explore other important applications, where ISS methods can help with optimization. Namely, we list a number of similar problems that can be solved in an analogous way, such as maximization of norms over a convex set, maximizing the Wigner-Yanase skew information, looking for the extremal eigenvalue of Hermitian matrices, and also for looking for states that violate maximally the Computable Cross Norm-Realignment (CCNR) criterion [30, 31].
@@
Our work is also related to recent efforts to solve various problems of numerical optimization in quantum metrology, including the combination of the tasks mentioned above, such as using an ISS for optimizing the quantum state and the unitary dynamics [23]. An ISS-based method has been used for an optimization over a general dynamics using its Choi-Jamiołkowski representation [32]. An ISS-based method has also been used for optimizing over adaptive strategies, when coherently probing several independent quantum channels [33]. An optimization over the probe state and the dynamics has been realized experimentally in a programmable quantum sensor realized in trapped cold ions [34]. For the identification of the optimal quantum metrological protocol in large systems, matrix product operators have been used [13]. Optimization methods, including ISS, has been considered in estimating the spectral density of a stochastic signal field [35]. Finally, numerical optimization in quantum metrology in various complex scenarios has been considered including multiparameter estimation [36, 37, 38, 39, 40, 41], and there are even methods that are based on semidefinite programmingx [42].
Variational methods used for the optimization of the probe state have already been mentioned. General optimization methods have also been applied for the optimization of the dynamics and other complex tasks. Optimization of the controls has been carried out using neural networks and reinforcement learning [43]. Machine learning approaches have been used in Bayesian parameter estimation [44]. The advantageous aspects of ISS methods to general optimization methods are discussed in Ref. [45].
@@
Our paper is organized as follows. In Sec. II, we review the basic notions of entanglement theory and quantum metrology. In Sec. III, we discuss, how our optimization problem can be viewed as a maximization over a convex set of Hamiltonians. We review the notion of metrological gain. In Sec. IV, we present various methods to optimize the quantum Fisher information over a local Hamiltonian. In Sec. V, we consider optimization problems in quantum information that can be solved in a similar way. For instance, we determine the bound entangled quantum states that maximally violate the CCNR criterion for entanglement detection [30, 31].
@@ II.1 Entanglement theory
In this section, we summarize basic notions of entanglement theory [1, 2]. A bipartite quantum state is called separable if it can be written as a mixture of product states as [55]
@@ II.2 Quantum metrology
Here is the Hamiltonian of the evolution, and is a parameter. Then, the metrological precision in estimating is bounded by the Cramér-Rao bound [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]
@@
the constant coefficients depend only on the eigenvalues of the probe state, and are given as
whenever the denominator is not zero. The matrix elements of the Hamiltonian in the basis of the eigenvectors of the density matrix are given as
@@
where the superscripts r and i denote the real and imaginary parts, respectively.
@@ III Basic notions for optimizing the metrological performance
First, not all Hamiltonians are easy to realize. In particular, local Hamiltonians are much easier to realize with a high precision than Hamiltonians with an interaction term. Local Hamiltonians appear most typically in quantum metrology forming the basis of linear interferometers. Thus, we will restrict the optimization for such Hamiltonians given as
@@
holds, where is a constant. Thus, it is easy to improve the metrological performance of any Hamiltonian by multiplying it by a constant factor. In order to obtain a meaningful optimization problem, we have to consider a formulation of such a problem where such a "trick" cannot be used to increase the quantum Fisher information to infinity. One option could be to divide the quantum Fisher information with the square of some norm or seminorm of the Hamiltonian, for which
@@
can be considered. In this paper, we argue that it is advantageous to compare the metrological performance of a state to the best performance of separable states. Therefore, we will divide the quantum Fisher information of the state by the maximum quantum Fisher information achievable by separable states given as for some as [74, 10]
where and denote the maximal and minimal eigenvalues, respectively.
@@ III.1 Seminorm for local Hamiltonians
Proof. We will now follow steps similar to those in the proof of Observation 1, showing that the three properties of seminorms are satisfied.
@@ III.2 Metrological gain
Note that is invariant under local unitaries thus
where the transformed state is defined as
Many entanglement measures, for instance, the Entanglement of Formation has the above property [1].
We consider local Hamiltonians, as we have discussed at the beginning of Sec. III. While we do not study it in this work, it would also be interesting to consider non-local Hamiltonians for a similar definition. However, computing the analogous quantity for that case is difficult, since calculating the maximum of an operator expectation value for separable states is a hard task [1]. Possibly, one can try to calculate the maximum for quantum states with a positive partial transpose (PPT) instead, which is an upper bound on the maximum for separable states, and it can be obtained via semidefinite programming (e. g., Refs. [80, 10]).
@@
The metrological gain is also defined in the multiparticle case. It has been shown that, for the metrological gain the inequality [24]
holds, where is the entanglement depth. It is assumed that the local Hamiltonians have identical maximal and minimal eigenvalues, which is fulfilled in typical many-particle systems [24]. Thus, the advantage of using the metrological gain to characterize the performance of the setup is that it is directly related to multipartite entanglement. Moreover, maximal metrological gain is possible only with a maximal entanglement depth.
We remark that there are other quantities to characterize the metrological performance. One can consider the difference of the quantum Fisher information and its maximum for separable states, and include the possibility of non-unitary dynamics from the point of view of resource theories. Considering the difference rather than the fraction makes the mathematical treatment easier. This way, all entangled states turn out to be a useful resource [25]. One can also consider the difference of the quantum Fisher information and a bound that depends on the state, where the difference is positive only for entangled states [26, 27, 28, 29]. The advantage of this approach is that the quantity can be computed analytically.
@@ IV.1 See-saw optimization based on a bilinear form
where contains the ’s and contains the ’s and the ’s. The matrix is a positive semidefinite symmetric matrix with real values. For the case of the quantum Fisher information, it is even a diagonal matrix. We will also consider other problems where it is a general positive semidefinite symmetric matrix. Minimizing such an expression over is a problem that can efficiently be solved. In general, minimizing a convex function under constraints is a task appearing often in quantum metrology. For instance, there are efficient ways to obtain the minimum of the quantum Fisher information over a given set of states [81, 82, 83]. On the other hand, maximizing the expression given in Eq. (54) over is difficult. A usual way of maximizing such an expression is as follows [84, 85, 86]. We replace the optimization over by a simultaneous optimization of the vectors and of a bilinear form as
@@
For completeness, let us prove Eq. (55) for the case of being a general positive semidefinite symmetric matrix. It is easy to see that the right-hand side is never smaller than the left-hand side. Now let us prove that the right-hand side is never larger than the left-hand side. Let us start from
Hence, follows that
where the equality is due to the fact that is symmetric. Then, at least one of the following two inequalities hold
After including an optimization over and , the equality given in Eq. (55) follows.
Based on the relation in Eq. (55), we can write the following.
@@
Based on Eq. (59), we can carry out the following algorithm. Initially, we choose randomly. Then, we maximize over the matrix while keeping fixed. Such an optimization can be carried out, for instance, by semidefinite programming [87] or by simple matrix algebraic calculations, see Appendix B. Then, we maximize over while keeping fixed. Then, we optimize again over , etc. This process is illustrated in Fig. 3. Note that we do not have to verify that the random initial Hamiltonian fulfills the constraints given in Eq. (49), since a random initial Hamiltonian and for any as initial Hamiltonian will lead to the same optimal (An alternative formulation of the optimization problem is in Appendix C. With similar methods, we also analyze the optimization over a quantum state rather than over the Hamiltonian Appendix D.)
@@
Since is convex in and if we maximize it over a convex set of Hamiltonians, then it will take the maximum at the boundary of the set. This optimization is a hard task, as we have discussed. We have to start from several random initial conditions, and only in some of the cases we will find a solution.
@@
where is the maximally entangled state and is the fraction of the white noise. In particular, we considered and . For such states, the optimal Hamiltonian is known analytically [23]. We also considered two copies of the state with and given as where the bipartite system consists of the parties and We also considered some bound entangled states that have been examined in the literature, and are often used to test various algorithms. Such states have a nontrivial structure, are highly mixed, but can be metrologically useful. We looked at the state based on an Unextendible Product Basis (UPB) [88, 89], the Horodecki state for [59], the metrologically useful bound entangled state in Ref. [10], and the private state given in Ref. [90, 11]. The results are shown in Table 1. The success probability of the method of Ref. [23] is similar, however, the formulation used in this paper makes it possible to design methods that provably find the optimum 111We started from a random initial Hermitian Hamiltonian, where the real and imaginary parts of the elements had a normal distribution with a zero mean and a unit variance, and they were independent from each other, apart from the requirement of the Hermiticity..
@@ IV.2 Quadratic programming with quadratic constraints
In the previous section, we considered methods that aim to maximize the quantum Fisher information over the Hamiltonian numerically. In practice, they are very efficient. Still, it is not certain that they find the global optimum and they might find a smaller value. In this section, we present methods that always find the global optimum. However, they work only for small systems. We also present simple approximations of such methods called relaxation. They find either the maximum or a value larger than that. The method given in Sec. IV.1 looking for the maximum with the see-saw and the methods discussed in this section can be used together. The gap between the two results give information about the location of the maximum. If the two methods find the same value, then it is the global maximum.
The basic idea of the method of moments can be understood as follows. According to the Shor relaxation [97, *[][.OriginallypublishedinSov.J.Comput.Syst.Sci.25, 1-11(1987).]Shor1987Quadratic], the maximization of Eq. (54) is replaced by the maximization of
@@
Note that Eq. (63) can easily be coded in semidefinite programming. We can also add further and further conditions on that lead to lower and lower values for the maximum. Similar relaxation techniques have been used in entanglement theory to detect entanglement, non-locality [99], to obtain the maximum of an operator for separable states [100], or even to determine the ground state energy in spin one-half Heisenberg models [101].
For our problem, we used the method of moments [102].
@@
After simple considerations, we can find that Eq. (66) is equivalent to Eq. (64) for all values. Thus, we do not need to carry out calculations for a range of values. A somewhat weaker constraint, which is computationally simpler is
We tested our methods for the same states we considered in Sec. IV.1. For the example, on a laptop computer, we could carry out calculations at level and Even level gave the exact result. For the examples, we could carry out calculations at level and and level gave the correct result. For the and examples, the level calculation gave already the correct result. The results are shown in Table 1. A detailed explanation of the method via a concrete example is in Appendix E.
@@ V.2 Maximization of the Wigner-Yanase skew information
Here, instead of we use since the latter is Hermitian. The expectation value of a non-Hermitian matrix could be complex. Then, the method finds the maximum based on the derivation presented in Eqs. (56), (57), and (58). Maximizing the Wigner-Yanase skew information can be used to lower bound the quantum Fisher information based on [108]
@@ V.3 Searching for the eigenstate with maximal/minimal eigenvalue
In this section, we discuss how to look for the largest eigenvalue of a positive semidefinite Hermitian matrix with a see-saw method similar to the one discussed in Sec. IV.1. Interestingly, the see-saw is guaranteed to converge in this case.
@@
where is a Hermitian matrix and the right-hand side of Eq. (72) can immediately be used to define a see-saw algorithm.
Let us now consider the Cauchy-Schwarz inequality
@@
Based on these considerations, the see-saw procedure is given as follows.
@@
With these steps, we realized the power method for finding the largest eigenvalue of a positive self-adjoint operator, which is a classic iterative algorithm in numerical linear algebra [109, 110]. The iteration is known to converge as long as the starting vector ( in our case) overlaps with the eigenvector corresponding to the maximal eigenvalue of . If is not positive semidefinite, then the method converges to the eigenvalue with the largest absolute value. Note that the problem could have also been solved with a one shot semidefinite program without iteration.